Ahmed is a football producer who is interested in knowing that at which level of production his profit would be maximized. If following is the production function determine at which quantity the profits would be maximized.


Y=10+10X+10X2

1
Expert's answer
2021-03-31T11:42:06-0400

The maximum of profits will be when the product function is maximum:


"\dfrac{dY}{dX}=0,""\dfrac{d}{dx}(10+10X-10X^2)=10-20X=0,""X=0.5"

Therefore, at "X=0.5" the profits would be maximized.